# Classification of thin Jordan schemes

### v1

@@ 4.1. Jordan algebraic automorphisms of coherent configurations
Let $\mathfrak{C}=(\Omega,C)$ be a coherent configuration, and let $\mathcal{A}$} be its adjacency algebra (considered as a Jordan algebra). Denote by $\Omega_{1},...,\Omega_{k}$ the fibers of $\mathfrak{C}$. Then the matrices $\underline{1_{\Omega_{i}}},i=1,...,k$ are the only minimal idempotents with respect to $\star$ and $\circ$. Therefore, every algebraic automorphism $\phi$ of $\mathfrak{C}$ permutes the relations $1_{\Omega_{i}},i=1,...,k$, that is, it acts on the set of indices 1,,k1,...,k. We abuse notation and write $\phi(1_{\Omega_{i}})=1_{\Omega_{\phi(i)}}$.
@@
Let $\mathfrak{C}=(\Omega,C)$ be the $2$-orbit configuration of a semi-regular subgroup $H\leq{\sf Sym}(\Omega)$. Then
@@
    $\mathfrak{C}\cong\mathsf{Inv}(\Omega_{1},H)\otimes\mathfrak{D}_{k}$ where $\Omega_{1}$ is an $H$-orbit, $k$ is the number of $H$-orbits on $\Omega$, and $\mathcal{D}_{k}$ is the discrete configuration on $k]:=\{1,\dots,k\}$;
@@
Proof of part (a). Let $\Omega_{1},...,\Omega_{k}$ be a complete set of $H$-orbits on $\Omega$. Clearly, $|\Omega_{1}|=...=|\Omega_{k}|=|H|$. Notice that a point set of $\mathsf{Inv}(\Omega_{1},H)\times\mathfrak{D}_{k}$ is $\Omega_{1}\times[k]$. Thus, to build a bijection $f:\Omega_{1}\times[k]\rightarrow\Omega$ let us choose arbitrary $\omega_{i}\in\Omega_{i},i=1,...,k$. Define $f(\omega,i)=h(\omega_{i})$ where $h\in H$ is a unique element satisfying $\omega=h(\omega_{1})$.
@@
Now, we check that $f$ maps basic relations of $(\Omega_{1},H)\otimes\mathfrak{D}_{k}$ onto basic relations of $\mathfrak{C}$. Since in both configurations all basic relations have the same cardinality, namely $|H|$, it is sufficient to prove that for any quadruple of points $(\alpha,i),(\beta,j),(\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime})\in\Omega_{1}\times[k]$ the following implication holds
$r((\alpha,i),(\beta,j))=r((\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime}))\Rightarrow r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime}))$
@@
It follows from the first equality that $i^{\prime}=i,j^{\prime}=j$ and $\alpha^{\prime}=h(\alpha),\beta^{\prime}=h(\beta)$ for a unique $h\in H$. Let $a,b\in H$ be elements satisfying $\alpha=a(\omega_{1}),\beta=b(\omega_{1})$. Then $\alpha^{\prime}=ha(\omega_{1}),\beta^{\prime}=hb(\omega_{1})$. Now we obtain $f(\alpha,i)=a(\omega_{i}),f(\beta,j)=b(\omega_{j})$ and $f(\alpha^{\prime},i)=ha(\omega_{i}),f(\beta^{\prime},j)=hb(\omega_{j})$. The latter equality implies $(f(\alpha^{\prime},i),f(\beta^{\prime},j))=(h(f(\alpha,i)),h(f(\beta,j)))$. Therefore, $r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime}))$.
Proof of part (b).
@@
First, we identify $\mathfrak{C}$ with $\mathsf{Inv}(\Omega_{1},H)\times\mathfrak{D}_{k}$. Under this identification a semi-regular action of $H$ on $\Omega_{1}\times[k]$ is given via $h(\omega,i)=(h(\omega),i)$. Denoting by $F$ the set of basic relations of $\mathsf{Inv}(\Omega_{1},H)$, we can now describe the basic relations of $\mathfrak{C}$ in the following form $f]_{i,j}=\{((\alpha,i),(\beta,j))\mid(\alpha,\beta)\in f\}$ where $f\in F$. In what follows we set $C:=\{[f]_{ij}\mid f\in F,i,j\in[k]\}$.
@@
Since $H$ acts regularly on $\Omega_{1}$, its 2-orbit scheme is thin, i.e. each $f\in F$ is a permutation on $\Omega_{1}$ and $F$ is a regular permutation subgroup of ${\sf Sym}(\Omega_{1})$. Notice that $F$ is a centralizer of $H$ in ${\sf Sym}(\Omega_{1})$ and $F\cong H$.
@@
Let us describe certain elements of the group $\mathsf{AAut}(\mathfrak{C})$. Given $v\in F^{k}$, define a permutation $\hat{v}\in{\sf Sym}(C)$ as follows $\hat{v}:[f]_{i,j}\mapsto[v_{i}^{-1}fv_{j}]_{i,j}$. It is easy to check that $\hat{v}$ is an algebraic automorphism of $\mathfrak{C}$. Another class of algebraic automorphisms comes from the symmetric group $S_{k}$: any permutation $\psi\in S_{k}$ permutes the basic relations in a following way $f]_{i,j}\mapsto[f]_{\psi(i),\psi(j)}$. We’ll denote these permutations as $\hat{\psi}$. The third type of algebraic automorphisms arrives from ${\sf Aut}(F)$, namely: any automorphism $\psi$ of $F$ permutes the basic relations as follows $f]_{i,j}\mapsto[\psi(f)]_{i,i}$. As before, we denote this induced permutation as $\hat{\psi}$. One can check that the three types of automorphisms described above generate a group isomorphic to $F^{k}\rtimes(S_{k}\times{\sf Aut}(F))$.
@@
Pick an arbitrary $\phi\in\mathsf{JAut}(\mathfrak{C})$. Then $\forall_{i\in[k]}\ \phi(I_{\Omega_{i}})=I_{\Omega_{{\varphi}(i)}}$ for a suitable ${\varphi}\in S_{k}$. Since $\hat{\varphi}\in\mathsf{AAut}(\mathfrak{C})$, $\hat{\varphi}^{-1}\phi$ is a Jordan automorphism of $\mathfrak{C}$ which fixes each $I_{\Omega_{i}},i\in[k]$. Thus in what follows we may assume that $\forall_{i\in[k]}\ \phi(I_{\Omega_{i}})=I_{\Omega_{i}}$. Consider the matrices $E_{i,j}:=\frac{1}{|H|}\underline{\Omega_{i}\times\Omega_{j}}$. The linear span ${\mathcal{A}}_{0}:=\langle E_{i,j}\rangle_{i,j\in[k]}$ is a two-sided ideal of ${\mathcal{A}}:=\mathbb{C}[C]$ isomorphic to $M_{k}(\mathbb{C})$. Since $E_{i,i}$ is unique non-zero solution of the equations $X\star I_{\Omega_{i}}=X,X\star X=X,X\circ X=\frac{1}{|H|}X$ and $\phi(I_{\Omega_{i}})=I_{\Omega_{i}}$, we obtain that $\phi(E_{i,i})=E_{i,i}$ for each $i\in[k]$. Given $i\neq j\in[k]$, the equations $E_{i,i}\star X=\frac{1}{2}X,E_{j,j}\star X=\frac{1}{2}X,X\circ X=\frac{1}{|H}X$ have twp non-zero solutions $E_{i,j}$ and $E_{j,i}$. Hence for any pair $i\neq j$ either $\phi(E_{i,j})=E_{j,i},\phi(E_{j,i})=E_{i,j}$ or $\phi(E_{i,j})=E_{i,j},\phi(E_{j,i})=E_{j,i}$. Let us call a pair switched in the first case and non-switched in the latter one.
@@
Our next step is to show that all pairs $i\neq j$ are either switched or non-switched simultaneously. This trivial if $k=2$. In the case of $k\geq 3$ it is sufficient to prove that for any three distinct elements $i,j,\ell\in[k]$ all pairs $\{i,j\},\{i,\ell\},\{j,\ell\}$ are non-switched provided that one of them is such.
@@
Assume, towards a contradiction, that $i,j$ is a non-switched pair, while $i,\ell$ is switched. Then $\phi(E_{i,j})=E_{i,j},\phi(E_{j,i})=E_{j,i}$ and $\phi(E_{i,\ell})=E_{\ell,i},\phi(E_{\ell,i})=E_{i,\ell}$. This implies
$E_{\ell,j}=2E_{i,j}\star E_{\ell,i}=2\phi(E_{i,j})\star\phi(E_{i,\ell})=2\phi(E_{i,j}\star E_{i,\ell})=0.$
A contradiction.
@@
The above claim implies that $\phi$ acts on ${\mathcal{A}}_{0}\cong M_{k}(\mathbb{C})$ either trivially or as a matrix transposition. Combining $\phi$ with t in the second case we may always assume that $\phi$ is trivial on ${\mathcal{A}}_{0}$, i.e. $\phi(E_{i,j})=E_{i,j}$ holds for all $i,j\in[k]$. The latter equality means that for any $i,j\in[k]$ the set of relations $\{[f]_{i,j}\}_{f\in F}$ is $\phi$ invariant, i.e. there exists a permutation $\phi_{ij}\in{\sf Sym}(F)$ such that $\phi([f]_{i,j})=[\phi_{ij}(f)]_{i,j}$.
@@
It follows from $f]_{ii}\star[h]_{ii}=[fh]_{ii}+[hf]_{ii}$ that $\phi_{ii}(fh)\in\{\phi_{ii}(f)\phi_{ii}(h),\phi_{ii}(h)\phi_{ii}(f)\}$. Thus $\phi_{ii}$ is a half-automorphism [13] of $F$. According to [13], $\phi_{ii}$ is either automorphism or anti-automorphism.
@@
For every pair $i\neq j$ and $f,h\in F$ we have $2[f]_{i,j}\star[h]_{j,i}=[fh]_{i,i}+[hf]_{j,j}$. Applying $\phi$ to both sides we obtain
@@
for all $f,h\in F$. Substituting $f=e$ into (14) we obtain
@@
It follows from $f_{ii}(e)=e$ that $\phi_{ij}(e)=\phi_{ji}(e)^{-1}$. Substituting $h=e$ into (14) we obtain $\phi_{ij}(f)=\phi_{ii}(f)\phi_{ji}(e)^{-1}$. Together with $\phi_{ji}(h)=\phi_{ij}(e)^{-1}\phi_{ii}(h)$ we obtain
@@
Now (14) implies that $\phi_{ii}(f)\phi_{ii}(h)=\phi_{ii}(fh)$, i.e. $\phi_{ii}$ is an automorphism of $F$ for every $i\in[k]$.
Since $\phi_{11}^{-1}\in{\sf Aut}(F)$, it induces an algebraic automorphism $\widehat{\phi_{11}^{-1}}\in\mathsf{AAut}(\mathfrak{C})$ which acts via $f]_{ij}\mapsto[\phi_{11}^{-1}(f)]_{ij}$. Replacing $\phi$ by $\phi\widehat{\phi_{11}^{-1}}$, we may assume that $\phi_{11}$ is trivial. This yields us $\phi_{jj}(h)=\phi_{j1}(e)^{-1}\phi_{11}(h)\phi_{1j}(e).$ Denoting $a_{j}:=\phi_{1j}(e)\in F$ we can write $\phi_{jj}(h)=h^{a_{j}}$ for each $j\in[k]$ (here $h^{a}$ stands for $a^{-1}ha$). Also $\phi_{1j}(f)=\phi_{11}(f)\phi_{j1}(e)^{-1}=fa_{j}$.
To finish the proof it is sufficient to show that $\phi=\hat{a}$ where $a:=(a_{1},...,a_{k})\in F^{k}$. The equality $\phi=\hat{a}$ is equivalent to $\phi([f]_{ij})=\hat{a}([f]_{ij})=[a_{i}^{-1}fa_{j}]_{ij}$.
It follows from (14) that $\phi_{ij}(f)=\phi_{ii}(f)\phi_{ji}(e)^{-1}=a_{i}^{-1}fa_{i}\phi_{ji}(e)^{-1}$. So, it remains to show that $\phi_{ji}(e)=a_{j}^{-1}a_{i}$. In the case of $i=j$ the equality follows from $\phi_{ii}\in{\sf Aut}(F)F)$. So, we may assume that $i\neq j$. It follows from $2[e]_{ji}=[e]_{j1}\star[e]_{1i}$ that $\phi_{ji}(e)=\phi_{j1}(e)\phi_{1i}(e)=\phi_{1j}(e)^{-1}\phi_{1i}(e)=a_{j}^{-1}a_{i}$, as desired. ∎

### v2

@@ 4.1. Jordan algebraic automorphisms of coherent configurations
Let $\mathfrak{C}=(\Omega,C)$ be a coherent configuration, and let $\mathcal{A}$} be its adjacency algebra which we consider as a Jordan algebra. Denote by $\Omega_{1},\dots,\Omega_{k}$ the fibers of $\mathfrak{C}$. Then the matrices $\underline{1_{\Omega_{i}}},i=1,\dots,k$ are the only minimal idempotents with respect to $\star$ and $\circ$. Therefore, every algebraic automorphism $\phi$ of $\mathfrak{C}$ permutes the relations $1_{\Omega_{i}},i=1,\dots,k$, which we may regard as a permutation of the indices $1,\dots,k$. Following this, we abuse notation by writing $\phi(1_{\Omega_{i}})=1_{\Omega_{\phi(i)}}$.
In what follows, if $r$ is a relation on $\Omega$ and $\Omega_{1}\subseteq\Omega$, then we write $r\mathord{\upharpoonright}_{\Omega_{1}}:=r\cap(\Omega_{1}\times\Omega_{1})$ to indicate the restriction of $r$ to $\Omega_{1}$. Moreover, if $R$ is a set of relations on $\Omega$ we write $R\mathord{\upharpoonright}_{\Omega_{1}}:=\{r\mathord{\upharpoonright}_{\Omega_{1}}\mid r\in R\}$.
@@
Let $\mathfrak{C}=(\Omega,C)$ be the $2$-orbit configuration of a semiregular subgroup $H\leq{\sf Sym}(\Omega)$. Then
@@
    $\mathfrak{C}\cong(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$ where $\Omega_{1}$ is an $H$-orbit, $k$ is the number of $H$-orbits on $\Omega$, and $\mathfrak{D}_{k}$ is the discrete configuration on $k]:=\{1,\dots,k\}$.
@@
Proof of (a). Let $\Omega_{1},\dots,\Omega_{k}$ be a complete set of $H$-orbits on $\Omega$. Clearly $|\Omega_{i}|=|H|$ for all $i\in[k]$. Note that $\Omega_{1}\times[k]$ is the point set of $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$. Thus, to build a bijection $f:\Omega_{1}\times[k]\rightarrow\Omega$ we choose arbitrary $\omega_{i}\in\Omega_{i},i\in[k]$. Define $f(\omega,i)=h(\omega_{i})$ where $h\in H$ is the unique element satisfying $\omega=h(\omega_{1})$.
@@
We now check that $f$ maps basic relations of $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$ onto basic relations of $\mathfrak{C}$. Since in both configurations all basic relations have the same cardinality, namely $|H|$, it is sufficient to prove that for any quadruple of points $(\alpha,i),(\beta,j),(\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime})\in\Omega_{1}\times[k]$ the following holds:
$r((\alpha,i),(\beta,j))=r((\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime}))\Rightarrow r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime})).$
@@
It follows from the first equality that $i^{\prime}=i$ and $j^{\prime}=j$. Also $\alpha^{\prime}=h(\alpha)$ and $\beta^{\prime}=h(\beta)$ where $h$ is the unique element of $H$ that maps $\alpha$ to $\alpha^{\prime}$. Let $a,b\in H$ be elements satisfying $\alpha=a(\omega_{1}),\beta=b(\omega_{1})$. Then $\alpha^{\prime}=ha(\omega_{1})$ and $\beta^{\prime}=hb(\omega_{1})$. Then we obtain $f(\alpha,i)=a(\omega_{i})$, $f(\beta,j)=b(\omega_{j})$ and $f(\alpha^{\prime},i)=ha(\omega_{i})$, $f(\beta^{\prime},j)=hb(\omega_{j})$. The latter equality implies $(f(\alpha^{\prime},i),f(\beta^{\prime},j))=(h(f(\alpha,i)),h(f(\beta,j)))$, whence we have $r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime}))$ proving (a).
Proof of (b). We first describe the group of isomorphisms $\mathsf{Iso}(\mathfrak{C})$ (referred to as color automorphisms in [10]). Since $H$ acts semiregularly on $\Omega$, $H$ is $2$-closed [18]. In this case $\mathsf{Iso}(\mathfrak{C})$ coincides with the normalizer of $H$ in the symmetric group ${\sf Sym}(\Omega)$.
@@
Under the identification $\mathfrak{C}$ with $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$. a semiregular action of $H$ on $\Omega_{1}\times[k]$ is given by $h(\omega,i)=(h(\omega),i)$. Denoting by $F$ the set of basic relations of $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})$, we can now describe the basic relations of $\mathfrak{C}$ in the following manner: $f]_{i,j}=\{((\alpha,i),(\beta,j))\mid(\alpha,\beta)\in f\}$ where $f\in F$. In what follows, we set $C:=\{[f]_{ij}\mid f\in F,i,j\in[k]\}$.
@@
Since $H$ acts regularly on $\Omega_{1}$, its 2-orbit scheme is thin, i.e. each $f\in F$ is a permutation on $\Omega_{1}$ and $F$ is a regular permutation subgroup of ${\sf Sym}(\Omega_{1})$. Note that $F$ is the centralizer of $H$ in ${\sf Sym}(\Omega_{1})$ and $F\cong H$.
@@
We next describe certain elements of the group $\mathsf{AAut}(\mathfrak{C})$. Given $v\in F^{k}$, we define a permutation $\hat{v}\in{\sf Sym}(C)$ by $\hat{v}:[f]_{i,j}\mapsto[v_{i}^{-1}fv_{j}]_{i,j}$. It is easy to check that $\hat{v}$ is an algebraic automorphism of $\mathfrak{C}$. Another class of algebraic automorphisms comes from the symmetric group $S_{k}={\sf Sym}([k])$, namely every permutation $\psi\in S_{k}$ permutes the basic relations in the following manner: $f]_{i,j}\mapsto[f]_{\psi(i),\psi(j)}$. We denote such induced permutations by $\hat{\psi}$.
A third class of algebraic automorphisms comes from ${\sf Aut}(F)$, namely $\psi\in{\sf Aut}(F)$ permutes the basic relations by $f]_{i,j}\mapsto[\psi(f)]_{i,j}$. In keeping with the above, we denote these induced permutations by $\hat{\psi}$. One can easily check that the three types of automorphisms just described generate a group isomorphic to $F^{k}\rtimes(S_{k}\times{\sf Aut}(F))$.
Pick an arbitrary $\phi\in\mathsf{JAut}(\mathfrak{C})$. Then $\forall_{i\in[k]}\ \phi(I_{\Omega_{i}})=I_{\Omega_{{\varphi}(i)}}$ for a suitably chosen ${\varphi}\in S_{k}$. Since $\hat{\varphi}\in\mathsf{AAut}(\mathfrak{C})$, we have that $\hat{\varphi}^{-1}\phi$ is a Jordan automorphism of $\mathfrak{C}$ that fixes each fiber $I_{\Omega_{i}},i\in[k]$. Thus, in what follows we may assume $\forall_{i\in[k]}\ \phi(I_{\Omega_{i}})=I_{\Omega_{i}}$.
@@
Consider now the matrices $E_{i,j}:=\frac{1}{|H|}\underline{\Omega_{i}\times\Omega_{j}}$. The linear span ${\mathcal{A}}_{0}:=\langle E_{i,j}\rangle_{i,j\in[k]}$ is a two-sided ideal of ${\mathcal{A}}:=\mathbb{C}[C]$ isomorphic to $M_{k}(\mathbb{C})$. Since $E_{i,i}$ is the unique nonzero solution to the equations $\star I_{\Omega_{i}}=X, XX$$\star X=X$, $X\circ X=\frac{1}{|H|}X$ and $\phi(I_{\Omega_{i}})=I_{\Omega_{i}}$, we obtain $\phi(E_{i,i})=E_{i,i}$ for each $i\in[k]$. Given $i\neq j\in[k]$, the equations $E_{i,i}\star X=\frac{1}{2}X$, $E_{j,j}\star X=\frac{1}{2}X$ and $X\circ X=\frac{1}{|H|}X$ have two nonzero solutions $E_{i,j}$ and $E_{j,i}$. Hence for any pair $i\neq j$, it is either the case that $\phi(E_{i,j})=E_{j,i},\phi(E_{j,i})=E_{i,j}$ or $\phi(E_{i,j})=E_{i,j},\phi(E_{j,i})=E_{j,i}$. Let us call a pair switched in the former case and non-switched in the latter case.
@@
Our next step is to show that if a single pair $\{i,j\}\subseteq[k]$ is switched, then all pairs in $[k]$ are switched. This is trivial if $k=2$. For $k\geq 3$, it suffices to prove that given any three distinct elements $i,j,\ell\in[k]$, if $\{i,j\}$ is switched then $\{i,\ell\}$ and $\{j,\ell\}$ are also switched.
@@
By way of contradiction, suppose $\{i,j\}$ is a switched pair and at least one of $\{i,\ell\}$ and $\{j,\ell\}$ is non-switched. First suppose $\{i,\ell\}$ is non-switched. Then $\phi(E_{i,j})=E_{j,i},\phi(E_{j,i})=E_{i,j}$ and $\phi(E_{i,\ell})=E_{i,\ell},\phi(E_{\ell,i})=E_{\ell,i}$. This implies
$E_{\ell,j}=2E_{i,j}\star E_{\ell,i}=2\phi(E_{j,i})\star\phi(E_{\ell,i})=2\phi(E_{j,i}\star E_{\ell,i})=0,$
a contradiction. Thus we are in the case where $\{i,j\}$ and $\{i,\ell\}$ are switched and $\{j,\ell\}$ is non-switched. But here, just as above, we reach the contradiction
$E_{i,j}=2E_{i,\ell}\star E_{\ell,j}=2\phi(E_{\ell,i})\star\phi(E_{\ell,j})=2\phi(E_{\ell,i}\star E_{\ell,j})=0.$
We conclude that every pair in $[k]$ is switched or every such pair is non-switched.
@@
It follows from above that $\phi$ acts on ${\mathcal{A}}_{0}\cong M_{k}(\mathbb{C})$ either trivially or via matrix transposition. Combining $\phi$ with t in the latter case, we may assume $\phi$ acts trivially on ${\mathcal{A}}_{0}$, i.e. $\phi(E_{i,j})=E_{i,j}$ for all $i,j\in[k]$. This implies the set of relations $\{[f]_{i,j}\}_{f\in F}$ is $\phi$-invariant for any $i,j\in[k]$, i.e. there exists a permutation $\phi_{ij}\in{\sf Sym}(F)$ such that $\phi([f]_{i,j})=[\phi_{ij}(f)]_{i,j}$.
@@
It follows from $f]_{ii}\star[h]_{ii}=[fh]_{ii}+[hf]_{ii}$ that $\phi_{ii}(fh)\in\{\phi_{ii}(f)\phi_{ii}(h),\phi_{ii}(h)\phi_{ii}(f)\}$. Thus $\phi_{ii}$ is a half-automorphism [15] of $F$. According to [15], $\phi_{ii}$ is either an automorphism or an anti-automorphism.
@@
For every pair $i\neq j$ and $f,h\in F$ we have $2[f]_{i,j}\star[h]_{j,i}=[fh]_{i,i}+[hf]_{j,j}$. Applying $\phi$ to both sides, we obtain
@@
for all $f,h\in F$. Substituting $f=e$ into (13) now gives
@@
It follows from $f_{ii}(e)=e$ that $\phi_{ij}(e)=\phi_{ji}(e)^{-1}$. Substituting $h=e$ into (13) we obtain $\phi_{ij}(f)=\phi_{ii}(f)\phi_{ji}(e)^{-1}$. Together with $\phi_{ji}(h)=\phi_{ij}(e)^{-1}\phi_{ii}(h)$ this yields
@@
It now follows from (13) that $\phi_{ii}(f)\phi_{ii}(h)=\phi_{ii}(fh)$, i.e. $\phi_{ii}$ is an automorphism of $F$ for every $i\in[k]$.
Since $\phi_{11}^{-1}\in{\sf Aut}(F)$, it induces an algebraic automorphism $\widehat{\phi_{11}^{-1}}\in\mathsf{AAut}(\mathfrak{C})$ which acts via $f]_{ij}\mapsto[\phi_{11}^{-1}(f)]_{ij}$. Replacing $\phi$ by $\phi\widehat{\phi_{11}^{-1}}$,we may assume that $\phi_{11}$ is trivial. This yields $\phi_{jj}(h)=\phi_{j1}(e)^{-1}\phi_{11}(h)\phi_{1j}(e).$ Setting $a_{j}:=\phi_{1j}(e)\in F$, we may write $\phi_{jj}(h)=a_{j}^{-1}ha_{j}$ for each $j\in[k]$. Also $\phi_{1j}(f)=\phi_{11}(f)\phi_{j1}(e)^{-1}=fa_{j}$.
To complete the proof, it suffices to show that $\phi=\hat{a}$ where $a:=(a_{1},\dots,a_{k})\in F^{k}$. The equality $\phi=\hat{a}$ is equivalent to $\phi([f]_{ij})=\hat{a}([f]_{ij})=[a_{i}^{-1}fa_{j}]_{ij}$.
It follows from (13) that $\phi_{ij}(f)=\phi_{ii}(f)\phi_{ji}(e)^{-1}=a_{i}^{-1}fa_{i}\phi_{ji}(e)^{-1}$. So it remains to show that $\phi_{ji}(e)=a_{j}^{-1}a_{i}$. If $i=j$, this equality follows from $\phi_{ii}\in{\sf Aut}(F)$. Hence we assume $i\neq j$. But now $2[e]_{ji}=[e]_{j1}\star[e]_{1i}$ implies $\phi_{ji}(e)=\phi_{j1}(e)\phi_{1i}(e)=\phi_{1j}(e)^{-1}\phi_{1i}(e)=a_{j}^{-1}a_{i}$ as desired. ∎